Write the slope-intercept form of the equation of the line perpendicular to the graph of
y = − 3 x − 7 that passes through 2
(3, –2).

Answer by mananth(16946)

1 y = -3 x -7
Divide by 1
y = -3 x + -7
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = -3

The slope of a line perpendicular to the above line will be the negative reciprocal 1/3
Because m1*m2 =-1
The slope of the required line will be 1/3

m= 1/3 ,point ( 3 , -2 )
Find b by plugging the values of m & the point in
y=mx+b
-2 = 1 + b
b= -3
m= 1/3
The required equation is y = 1/ 3 x -3
m.ananth@hotmail.ca